IOK 12K 3 WITH A tt 8,89 100040 200046 250 PLF tot 20find the support reactions
Answer #1


solution. Gok 2 x 30 = w= 2*1f7 lok 12 K G+EMA=0; B. A 17.5 8 8 TOXO - GOX1705 + By x x 38- 12x46=0 301 Ay fig By GOX 17.5+

Please Rate

Thank you !!!!!!!!